a meeting of consuls was attended by 4 Americans and 4 Germans. If three consuls were selected at random, construct the probability distribution of the random variable G representing the number of Germans.






1
Expert's answer
2022-02-16T07:55:32-0500

Let "A" represent the number of Americans, "G" represent the number of Germans.

We can select three consuls from "4+4=8" consuls in "\dbinom{8}{3}=56" ways.

The possible values of random variable "G" representing the number of Germans are "0,1,2,3."


"P(G=0)=\dfrac{\dbinom{4}{0}\dbinom{4}{3-0}}{\dbinom{8}{3}}=\dfrac{1(4)}{56}=\dfrac{1}{14}"

"P(G=1)=\dfrac{\dbinom{4}{1}\dbinom{4}{3-1}}{\dbinom{8}{3}}=\dfrac{4(6)}{56}=\dfrac{3}{7}"

"P(G=2)=\dfrac{\dbinom{4}{2}\dbinom{4}{3-2}}{\dbinom{8}{3}}=\dfrac{6(4)}{56}=\dfrac{3}{7}"

"P(G=3)=\dfrac{\dbinom{4}{3}\dbinom{4}{3-3}}{\dbinom{8}{3}}=\dfrac{4(1)}{56}=\dfrac{1}{14}"

The probability distribution of the random variable G representing the number of Germans is

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 g & 0 & 1 & 2 & 3 \\ \hline
 p(g) & 1/14 & 3/7 & 3/7 & 1/14 
\end{array}"